# JanusCoder: Towards a Foundational Visual-Programmatic Interface for Code Intelligence

### v1

@@ 2 Related Works
LLMs have been widely explored for text-centric code generation of visual interfaces, including data visualizations (Yang et al., 2024), web pages (Chen et al., 2025d), and interactive UIs (Chen et al., 2025a). Early efforts focused on Python libraries (e.g., Matplotlib, Seaborn) for producing figures in scientific workflows (Zhang et al., 2024b; Sun et al., 2025b). Later work extended to chart generation and editing (Zhao et al., 2025a), and to mapping NL instructions into web-based artifacts (Zhang et al., 2025) or structured UI interactions (Cheng et al., 2024; Sun et al., 2024b). Overall, these approaches highlight the potential of LLMs to author executable visual content, though they remain constrained to text-driven inputs.
@@
Another line of work emphasizes multimodal inputs (vision-centric), where models interpret visual information to produce or reason about symbolic code (Hu et al., 2024; Jiang et al., 2025). Representative efforts include chart understanding, which evaluates the extraction of structured knowledge from plots (Masry et al., 2022; Zhang et al., 2024a), and chart-to-code generation, which requires reproducing scientific plots from images with captions or instructions (Zhao et al., 2025b; Xia et al., 2025; Wu et al., 2025). Beyond charts, studies extend to theorem visualization (Ku et al., 2025), multimodal algorithmic problem solving (Li et al., 2024), and structured vector graphics such as SVGs (Yang et al., 2025c; Nishina & Matsui, 2024). While these works demonstrate progress, they largely target isolated domains and modalities. In contrast, we move beyond these constraints by unifying diverse domains and modalities across charts, web UIs, animations, symbolic computation, and more, taking a leap forward in advancing multimodal code intelligence.
@@ 3 Method
To empower models for multimodal code intelligence, we propose a versatile data toolkit that incorporates model interactions (Sun et al., 2023b) and compiler feedback to tackle multifaceted demands. In contrast to prior data approaches, which often suffer from a lack of instruction diversity, scarcity in specialized domains, and insufficient validation for visual-code alignment, our pipeline establishes a principled workflow. As shown in Figure 2: (1) Data Sourcing, where raw assets are collected and categorized; (2) Data Synthesis & Curation, where new instruction-code pairs is generated and refined through a multi-strategy engine; and (3) Quality Control, which ensures data fidelity through automated validation and LLM/VLM judging.
@@ 4.2 Evaluation
The overall score is defined as $s_{\text{exec}}\cdot\big(s_{\text{sim}}+s_{\text{align}}+s_{\text{faith}}\big)$. This ensures that only executable code is considered for further evaluation, while successful generations are rewarded for syntactic fidelity, semantic alignment, and perceptual faithfulness. More details of DTVBench are in Appendix C.
@@ 5.1 Experimental Settings
For model construction, we use Qwen3-{8B,14B} (Yang et al., 2025a) as the backbones of JanusCoder, and Qwen2.5-VL-7B-Instruct (Bai et al., 2025) together with InternVL3.5-8B (Wang et al., 2025b) as the backbones of JanusCoderV. In the analysis part, we additionally include Qwen3-4B, Qwen2.5-Coder-7B-Instruct (Hui et al., 2024), and InternVL3.5-4B for further comparison. Model details are provided in Appendix D.
@@ 5.2 Benchmarking
We also evaluate on BigCodeBench (Zhuo et al., 2025) and LiveCodeBench (Jain et al., 2025) to highlight its capability in following complex instructions and algorithmic capability.
@@ 5.4 Main Results: Multimodal Tasks
We evaluate JanusCoderV on ChartMimic (Yang et al., 2025b), JanusCoderV achieves strong results on both high- and low-level metrics, consistently outperforming baselines and substantially surpassing GPT-4o. As a unified model, it also outperforms recently released specialized chart-to-code MLLMs (Xia et al., 2025; Zhao et al., 2025b), highlighting the effectiveness of leveraging cross-task data synergy. Detailed comparisons are provided in Appendix F.
@@
Due to space limitations, the detailed metrics for the results on all the aforementioned benchmarks are presented in Appendix E.
@@ 6.1 Ablation Studies
To validate the cross-domain and cross-modal synergies proposed in Section 3.3, we conduct ablation studies by selectively removing specific categories of data within JanusCode-800K. The results support our claim, showing that data from non-target domains, even when cross-modal, can provide transferable coding capabilities (Sun et al., 2023a) on specialized visual tasks (e.g., text-centric data contributing to multimodal coding scenarios). This provides useful guidance for the research community, suggesting that performance in data-scarce scenarios such as animations and artifacts can be improved by incorporating data from related, more abundant sources.
@@ 6.2 Effect of Backbones
This confirms the soundness of our data design and can empower diverse backbones to become more generalist models for multimodal code intelligence. More experiments on different backbones are available in Appendix F.2.
@@ 6.3 General Coding Capabilities
JanusCoder demonstrates superior general coding capabilities that surpass even specialist approaches. As shown in Figure 5, it achieves strong performance on general benchmarks while also outperforming specialist models like VisCoder (Ni et al., 2025) in their own target visualization domain. Furthermore, it outperforms GPT-4o in both scenarios, which further demonstrates our model’s balanced capabilities. More comparisons are provided in Appendix F.1.
@@ Appendix B Data Collection Details
Appendix C DTVBench Details
@@
Appendix D Training Details
@@
Appendix E Detailed Experimental Results
E.1 Detailed results on PandasPlotBench
@@
E.2 Detailed results on ChartMimic
@@
E.3 Detailed results on DesignBench
@@
E.4 Detailed results on WebCode2M
@@
E.5 Detailed Results on InteractScience
@@
E.6 Detailed results on ArtifactsBench
@@
Appendix F Detailed Analysis and Comparisons
F.1 General Coding Capabilities
@@
F.2 Experiments on Different Backbones
@@
Appendix G Case Studies

### v2

@@ Abstract
11footnotetext: *Equal contribution  Corresponding author Project lead
@@ 2 Related Works
LLMs have been widely explored for text-centric code generation of visual interfaces, including data visualizations (Yang et al., 2024; Zhuo et al., 2025b), web pages (Chen et al., 2025d), and interactive UIs (Chen et al., 2025a). Early efforts focused on Python libraries (e.g., Matplotlib, Seaborn) for producing figures in scientific workflows (Zhang et al., 2024b; Sun et al., 2025b). Later work extended to chart generation and editing (Zhao et al., 2025a), and to mapping NL instructions into web-based artifacts (Zhang et al., 2025) or structured UI interactions (Cheng et al., 2024; Sun et al., 2024b). Overall, these approaches highlight the potential of LLMs to author executable visual content, though they remain constrained to text-driven inputs.
@@
Another line of work emphasizes multimodal inputs (vision-centric), where models interpret visual information to produce or reason about symbolic code (Hu et al., 2024; Jiang et al., 2025). Representative efforts include chart understanding, which evaluates the extraction of structured knowledge from plots (Masry et al., 2022; Zhang et al., 2024a), and chart-to-code generation, which requires reproducing scientific plots from images with captions or instructions (Zhao et al., 2025b; Xia et al., 2025; Wu et al., 2025). Beyond charts, studies extend to theorem visualization (Ku et al., 2025), multimodal algorithmic problem solving (Li et al., 2024), and structured vector graphics such as SVGs (Yang et al., 2025c; Nishina and Matsui, 2024). While these works demonstrate progress, they largely target isolated domains and modalities. In contrast, we move beyond these constraints by unifying diverse domains and modalities across charts, web UIs, animations, symbolic computation, and more, taking a leap forward in advancing multimodal code intelligence.
@@ 3 Method
To empower models for multimodal code intelligence, we propose a versatile data toolkit that incorporates model interactions (Sun et al., 2024c) and compiler feedback to tackle multifaceted demands. In contrast to prior data approaches, which often suffer from a lack of instruction diversity, scarcity in specialized domains, and insufficient validation for visual-code alignment, our pipeline establishes a principled workflow. As shown in Figure 2: (1) Data Sourcing, where raw assets are collected and categorized; (2) Data Synthesis & Curation, where new instruction-code pairs is generated and refined through a multi-strategy engine; and (3) Quality Control, which ensures data fidelity through automated validation and LLM/VLM judging.
@@ 4.2 Evaluation
The overall score is defined as $s_{\text{exec}}\cdot\big(s_{\text{sim}}+s_{\text{align}}+s_{\text{faith}}\big)$. This ensures that only executable code is considered for further evaluation, while successful generations are rewarded for syntactic fidelity, semantic alignment, and perceptual faithfulness. More details of DTVBench are in Appendix D.
@@ 5.1 Experimental Settings
For model construction, we use Qwen3-{8B,14B} (Yang et al., 2025a) as the backbones of JanusCoder, and Qwen2.5-VL-7B-Instruct (Bai et al., 2025) together with InternVL3.5-8B (Wang et al., 2025b) as the backbones of JanusCoderV. In the analysis part, we additionally include Qwen3-4B, Qwen2.5-Coder-7B-Instruct (Hui et al., 2024), and InternVL3.5-4B for further comparison. Model details are provided in Appendix E.
@@ 5.2 Benchmarking
We also evaluate on BigCodeBench (Zhuo et al., 2025a) and LiveCodeBench (Jain et al., 2025) to highlight its capability in following complex instructions and algorithmic capability.
@@ 5.4 Main Results: Multimodal Tasks
We evaluate JanusCoderV on ChartMimic (Yang et al., 2025b), JanusCoderV achieves strong results on both high- and low-level metrics, consistently outperforming baselines and substantially surpassing GPT-4o. As a unified model, it also outperforms recently released specialized chart-to-code MLLMs (Xia et al., 2025; Zhao et al., 2025b), highlighting the effectiveness of leveraging cross-task data synergy. Detailed comparisons are provided in Appendix G.
@@
Due to space limitations, the detailed metrics for the results on all the aforementioned benchmarks are presented in Appendix F.
@@ 6.1 Ablation Studies
To validate the cross-domain and cross-modal synergies proposed in Section 3.3, we conduct ablation studies by selectively removing specific categories of data within JanusCode-800K. The results support our claim, showing that data from non-target domains, even when cross-modal, can provide transferable coding capabilities (Sun et al., 2023) on specialized visual tasks (e.g., text-centric data contributing to multimodal coding scenarios). This provides useful guidance for the research community, suggesting that performance in data-scarce scenarios such as animations and artifacts can be improved by incorporating data from related, more abundant sources.
@@ 6.2 Effect of Backbones
This confirms the soundness of our data design and can empower diverse backbones to become more generalist models for multimodal code intelligence. More experiments on different backbones are available in Appendix G.2.
@@ 6.3 General Coding Capabilities
JanusCoder demonstrates superior general coding capabilities that surpass even specialist approaches. As shown in Figure 5, it achieves strong performance on general benchmarks while also outperforming specialist models like VisCoder (Ni et al., 2025) in their own target visualization domain. Furthermore, it outperforms GPT-4o in both scenarios, which further demonstrates our model’s balanced capabilities. More comparisons are provided in Appendix G.1.
@@ 7 Conclusion
Acknowledgement
We thank Fangzhi Xu and Zhangyue Yin for their valuable feedback on improving the manuscript’s presentation. We also express our gratitude to Zichen Ding for the assistance in addressing challenges encountered during the fine-tuning of InternVL 3.5. This research is supported by Shanghai Artificial Intelligence Laboratory and WYNG Foundation (HKU 25AG100407), with sincere appreciation for their support.
Author Contributions
The authors contributed to this work in the following ways: Qiushi Sun led the project, proposing the first feasible version of building a visual-programmatic interface and planning the overall direction. He implemented the initial data pipelines (Python, Matlab, R, Manim, and other PLs/libraries), proposed the data mixture strategy, wrote the main paper, and created the figures. Jingying Gong was responsible for overall model training (unimodal and multimodal) and evaluation of Python-related tasks, and resolved multiple engineering challenges in large-scale code data rollout. Yang Liu implemented the data pipeline and evaluation for HTML-related data, participated in multimodal training, and assisted in formulating the data mixture strategy. Qiaosheng Chen curated the seed data for interactive content and was responsible for the data pipeline and evaluation of demo-related data. Qiushi Sun, Jingyang Gong, Yang Liu, and Qiaosheng Chen all participated in paper reviewing, refinement, and providing demo cases. Lei Li assisted in discussions on code data synthesis. Kai Chen reviewed the submitted manuscript, provided detailed revision suggestions, and discussed strategies for better data utilization. Qipeng Guo oversaw the design of the main data strategy and provided substantial assistance in data pipelines, evaluation, and base model selection, and also offered valuable feedback during the paper rebuttal phase. Ben Kao offered advice on paper writing and impact maximization. Fei Yuan participated in discussions.
@@ Appendix B Data Collection Details
Appendix C Reward Modeling Details
To ensure the quality of our synthetic data, we implement an automated reward mechanism that systematically filters generated code samples and their corresponding multimodal outputs.
C.1 Rewarding Text-centric Data
For text-centric data lacking visual outputs (e.g., algorithmic implementations or symbolic computations), we utilize an LLM as a judge. We deploy Qwen-2.5-72B-Instruct to evaluate samples across dimensions, including instruction alignment, logical coherence, and code robustness. Scoring adheres to the criteria detailed in the Reward Prompt - Artifacts (Appendix I), where the model provides a rating from 1 to 5. Only samples achieving a score of 5 are retained to ensure high-fidelity instruction following in the final dataset.
C.2 Rewarding Vision-centric Data
For vision-centric data (e.g., Python visualizations, WebUI layouts, and animations), we employ Qwen2.5-VL-72B-Instruct to verify the alignment between instructions, code, and rendered visual outputs.
The rating for vision-centric data also follows a 5-point standard, but unlike text-centric data, the VLM-based reward function requires a more rigorous verification process:
  • Multimodal Alignment: The VLM processes the raw instruction $I$, the code $C$, and the rendered output $V$ to assess whether visual elements—such as labels, color schemes, and layout—strictly adhere to the input instructions.
  • Evaluation of Editing Tasks: For visual editing tasks (Viscode Edit), we apply a binary filtering strategy. As specified in the Reward Prompt - Viscode Edit (Appendix I), samples are marked as valid only if they achieve a minimum threshold ($\geq 3$) across key dimensions, including Task Completion and Visual Clarity.
By leveraging this reward mechanism, we filter the raw synthetic pool to construct the final JanusCode-800K corpus.
Appendix D DTVBench Details
@@
Appendix E Training Details
@@
Appendix F Detailed Experimental Results
F.1 Detailed results on PandasPlotBench
@@
F.2 Detailed results on ChartMimic
@@
F.3 Detailed results on DesignBench
@@
F.4 Detailed results on WebCode2M
@@
F.5 Detailed Results on InteractScience
@@
F.6 Detailed results on ArtifactsBench
@@
Appendix G Detailed Analysis and Comparisons
G.1 General Coding Capabilities
@@
G.2 Experiments on Different Backbones
@@
G.3 Cross-Model Adjudication
To mitigate potential evaluation circularity and bias from relying on a single judge backbone in our reward modeling pipeline, we conducted a cross-model adjudication study. We randomly sampled 30K unfiltered data samples (covering Python and WebUI artifacts) and evaluated them using three distinct frontier models: Qwen2.5-VL-72B-Instruct, Gemini-2.5-Pro, and GPT-5.
Table 14: Cross-Model Adjudication
Qwen2.5-VL (Mean) Qwen2.5-VL (Var) Gemini2.5-Pro (Mean) Gemini2.5-Pro (Var) GPT-5 (Mean) GPT-5 (Var) Krippendorff’s $\alpha$
Python 4.59 0.41 4.36 0.58 4.41 0.53 0.76
Artifacts 4.28 0.47 4.05 0.61 4.16 0.52 0.71
The evaluation yielded a Krippendorff’s $\alpha$ above 0.70 across both domains, indicating strong inter-annotator agreement among different model families. This confirms that our reward scores are robust, objective, and generalizable indicators of data quality.
G.4 Human Evaluation on Subjective Tasks
To further validate the subjective visual quality of our generated artifacts, which is often difficult to capture fully via automated metrics, we conducted a human evaluation on a subset of 100 generated samples (comprising Manim animations and InteractScience WebUIs). Three annotators with college-level backgrounds evaluated the outputs on a 1-5 scale based on visual fidelity and instruction alignment.
Table 15: Human Evaluation on Subjective Tasks
Model Manim Code2Video Model InteractScience WebUI
Qwen3-8b 2.1 Qwen-2.5-VL-7b 0.92
Qwen3-14b 2.64 InternVL-3.5-4B 1.67
Qwen-2.5-coder-7b 2.48 InternVL-3.5-8B 1.57
Qwen-2.5-coder-32b 2.82 GPT-4o 2.78
JanusCoder-14B 3.14 JanusCoderV-7B 3.19
As shown in Table 15, the JanusCoder series consistently outperforms the baseline models, correlating strongly with our automated VLM-based evaluations and confirming its superior ability to generate perceptually high-quality visual content.
Appendix H Case Studies